How do you factor (x^3+8)?
Answer by stanbon(48571)

factor (x^3+8)
---------------
Form: a^3+b^3 = (a+b)(a^2-ab+b^2)
Your Problem:
x^3+2^2 = (x+2)(x^2-2x+4)
==============
Cheers,
Stan H.